# Predicting wide binaries and deviations from standard gravity using machine learning algorithms

### v1

@@ IV.1 Deviations from standard gravity
In context of machine learning, we treat the deviations from standard gravity as an anomaly detection problem for which we require a dataset to train the ML model. This dataset is another catalogue generated by Chae 2023 ([20]), based on El-Badry’s work.
This catalogue has the binary pairs that are within $80pc$ distance and $200pc$ distance separately from the point of observation. However, for the larger distance, measurement uncertainties increase, while for the smaller distance, the dataset is smaller in number. Moreover, all the binary pairs are of ”MSMS” class. An ”MS” type star can be a giant, a subgiant, a pre-main sequence star or a brown dwarf. Some other constraints of this catalogue are:
@@
Another feature of this catalogue which is of utmost importance in determining the parameter for the training of our ML model is the mass-magnitude relation. It associates the luminosity of the star with its mass using the spectroscopic and astrometric data like parallax, G-magnitude, photon flux and extinction coefficient. Pecaut and Mamjek 2013 [49] provides the mass-magnitude relation for the Chae’s sources along with masses, several colours and magnitudes for a wide range of spectral types in various wave bands in a tabulated format. However, [49] has made use of the GAIA DR2 observations, whereas, in this work we are dealing with the eDR3 quantities. Hence, the other band magnitudes from DR2 to the G-magnitude in eDR3 data have been transformed [20]. The first way is to transform the V-band magnitudes([50])
$G-V=-0.01597-0.02809X_{VI}-0.2483X^{2}_{VI}+0.03656X^{3}_{VI}-0.002939X^{4}_{VI}$ (5)
where $X_{VI}\equiv V-I_{C}$. The second option is to use the 2MASS J-band magnitudes
$G-J=0.01798+1.389X_{BR}-0.09338X^{2}_{BR}$ (6)
where $X_{BR}\equiv BP-RP$.
@@
The polynomial with the coefficients in $V$ and $J$ band is given as
$log_{10}(\frac{M_{*}}{M_{o}})=\sum_{i=0}^{10}a_{i}(M_{G})^{i}$ (7)
@@
We consider the stars for analysis whose masses are $M_{*}<1.2M_{o}$, where $M_{o}$ is solar mass. However, the masses of other objects outside the scope of this work can be obtained.
Since this work aims at training an ML model only, the resultant table from Chae ([20]), which is hosted at Chae Scripts [51] has been directly accessed in this work. Another way of finding the mass magnitude directly using the G-band magnitudes from the eDR3 data is by using the mass-luminosity relation given by Chevalier 2023 [52]. The Gaia DR3 data provides the G-band magnitudes for all the binary star systems (Gaia Collaboration 2023a [42]. The Gaia dust extinction law is useful in this case to determine the masses[53]. However, the outcome polynomial is applicable only for a limited range of stellar mass $(0.12<M<0.77)$. As a conclusion, we draw that this method is applicable for our catalogue’s range $(M\leq 1.2M_{o}$) if we extrapolate the polynomial or through some simulating methods. Thus, this deviates our work from the objective of training an ML model, and we follow the older method.
@@ IV.2 Labelling deviations
In the training of ML model, we intentionally eliminated the positional information so as to avoid any spatial bias. This elimination also hints at proposing a new parameter for anomaly detection, besides cutting down the computational cost.
In literature (Pittordis [18], Banik et al.[9]), the deviation from Newtonian gravity has been inferred using the $\tilde{v}$ parameter. For the training of our ML model to identify anomalous systems, we employ this parameter as our primary input. $\tilde{v}$ is given as
@@
$v_{p}=4.7404\times 10^{-3}km/s\times\Delta\mu\times d$ (9)
$d$ is the distance in pc to the binary system and $\Delta\mu$ is the plane-of-sky proper motion given in section III.1.
$M_{tot}$ is the total mass of the system and $s$ is the separation between the two stars.
While other authors have considered the projected components of the velocity and separation, we make use of the 2D plane-of-sky velocity($v_{p}$) and the separation(s) estimated while making the catalogue itself. This is due to the fact that we don’t want to have a precise identification of a binary system that is deviating, but rather develop an anomaly detection system by a different approach. So to some extent, the uncertainties occurring are tolerable.
More recently, Chae([20]) has defined a new parameter using ’MCMC’ simulations to show deviations and discarded the idea of $\tilde{v}$ parameter, as the advent of projection effects makes it difficult to get correct $\tilde{v}$ distribution.
For our work, however, we follow the method of Pittordis([18]) of $\tilde{v}$ parameter because (i) for training an ML model, the computational cost is optimised and once the model is trained, the anomalies beyond our limits can also be calculated. (ii) Moreover, our model is trained on the MC realised results for Newtonian gravity(Chae[20]), thus we are aiming at more accurate results through the model.
To summarise, the MCMC procedure, being computationally expensive, is discarded. However, its results are used for the training of the ML model, which analyses the $\tilde{v}$ parameters to give anomalies.
As a final step to this work, we found the $\tilde{v}$ for the Newtonian simulated data points (extracted from a site mentioned in [20]) and the actual data points (of catalogue). We define a new parameter
$\begin{split}\Delta\tilde{v}=\tilde{v}_{New}-\tilde{v}_{obs}\\ \tilde{v}_{New}=\frac{v_{New}}{\sqrt{GM/s}}\\ \tilde{v}_{obs}=\frac{v_{p}}{\sqrt{GM/s}}\end{split}$ (10)
@@
$\Delta_{dev}=\tilde{v}_{actual}-\tilde{v}_{newtonian}$ (15)
@@
Start ProcessLoad Filtered Chae DatasetLoad Chae’s Newtonian Simulation DatasetCalculate $v_{p}$ and $v_{N}$alculate $\tilde{v}$alculate $v_{p}$ and $v_{N}$alculate $\tilde{v}$alculate the $\Delta_{dev}$ ParameterRemove Deviations Inside ±$\pm 3\sigma$Mark Deviations in the Chae DatasetEnd Process

### v2

@@ IV.1 Deviations from standard gravity
In context of machine learning, we treat the deviations from standard gravity as an anomaly detection problem for which we require a dataset to train the ML model. We use the catalogue generated by Chae as our training/testing/validation dataset [20]. This catalogue has the binary pairs that are within $80pc$ distance and $200pc$ distance separately from the point of observation. However, for the larger distance, measurement uncertainties increase, while for the smaller distance, the dataset is smaller in number. Moreover, all the binary pairs are of ”MSMS” class. An ”MS” type star can be a giant, a subgiant, a pre-main sequence star or a brown dwarf. Some other constraints of this catalogue are:
@@
$log_{10}(\frac{M_{*}}{M_{o}})=\sum_{i=0}^{10}a_{i}(M_{G})^{i},$ (5)
where the coefficients $a_{i}$ corresponding to both $V$ and $J$ bands as obtained in Ref. [20] are given in table 3. The mass is expressed in solar mass units $M_{o}$ and $M_{*}<1.2M_{o}$ for our purposes.
@@
Table 3: Coefficients of the polynomial expression for the mass-magnitude relation as given by Chae [20]. The coefficients are applicable for both $M_{V}$ and $M_{J}$ based G-band magnitudes. The mass is expressed in solar mass units $M_{o}$.
@@ IV.2 Labelling deviations
In the training of ML model, we intentionally eliminated the positional information so as to avoid any spatial bias. In literature [18, 9], the deviation from Newtonian gravity has been inferred using the $\tilde{v}$ parameter. For the training of our ML model to identify anomalous systems, we employ this parameter as our primary input. $\tilde{v}$ is given as
@@
$v_{p}=4.7404\times 10^{-3}km/s\times\Delta\mu_{rel}\times d$ (7)
$d$ is the distance in pc to the binary system and $\Delta\mu_{rel}$ is the plane-of-sky proper motion given in section III.1. $M_{tot}$ is the total mass of the system and $s$ is the separation between the two stars.
While the usual practice is to consider the projected components of the velocity and separation, we make use of the 2D plane-of-sky velocity($v_{p}$) and the separation(s) estimated while making the catalogue itself that requires no additional computational steps. We note that $\tilde{v}$ or $v_{p}$ are not accurate measures of deviations. In fact, Chae [20] has defined new parameters like $\delta_{newt-obs}$ which are more accurate but require using MCMC simulations. For our work, however, we use the $\tilde{v}$ parameter defined in Ref. [18] because (i) for training an ML model, the computational cost is optimised and once the model is trained, the anomalies beyond our limits can also be calculated. (ii) our model is trained on the MC realised results for Newtonian gravity [20], thus we are aiming at more accurate results through the model. In summary, the MCMC procedure, being computationally expensive, is discarded. However, its results are used for the training of the ML model, which analyses the $\tilde{v}$ parameters to give anomalies.
@@
$\Delta_{dev}=\tilde{v}_{obs}-\tilde{v}_{newtonian}$ (12)
@@
Start ProcessLoad Filtered Chae DatasetLoad Chae’s Newtonian Simulation DatasetCalculate $v_{p}$ and $v_{N}$alculate $\tilde{v}_{obs}$alculate $v_{p}$ and $v_{N}$alculate $\tilde{v}_{newtonian}$alculate the $\Delta_{dev}$ ParameterRemove Deviations Inside ±$\pm 3\sigma$Mark Deviations in the Chae DatasetEnd Process